OpenStudy (anonymous):

Write the linear function for the line with slope 5/4 that passes through the point (-4, 1) and graph the function. Please show all of your work Write the linear function for the line with slope 5/4 that passes through the point (-4, 1) and graph the function. Please show all of your work @Mathematics

OpenStudy (anonymous):

The equation of a line in slope-intercept form is y=mx+b You know that the slope is 5/4, so you can just rewrite the equation as y=(5/4)x+ b Then you just have to find b. To find b, plug in the point you are given, put -4 for x and 1 for y and solve for b. To graph the function, you can plug in numbers and just connect the dots.

OpenStudy (anonymous):

I dont understand how to solve for b

OpenStudy (anonymous):

The point (-4,1) represents (x,y). To find the linear function, all you have to do is plug in the x and the y into your equation. So when you have y=(5/4)x+b you can put 1=(5/4)(-4)+b. After you simplify this, move the number to one side and make it equal to b. That b is your b for the equation.

OpenStudy (anonymous):

ok, so b = -5 ? Is this correct?

OpenStudy (anonymous):

Close, (5/4) times (-4) is -5, but you can't forget about the 1 you have on the other side of your equation. So at that step, you have 1=-5+b To solve for b, you have to move -5 over to the other side, so you get 6=b

OpenStudy (anonymous):

ok so how would I graph this then?

OpenStudy (anonymous):

IM so confused

OpenStudy (anonymous):

Think about your equation as a machine. You put in an y and it gives you an x, or you can put in an x and it gives you a y. For example, if I say x=1 (just a random number, any will work) Y=(5/4)1+6 After you simplify you get y=29/4 So a point on your graph is (1,(29/4)) However, it's easier to use whole numbers. You can use the point you already have (-4,1) and just go from there. Plot (-4,1) and make another point 5 points above (-4,1) and 4 points to the right of it

OpenStudy (anonymous):

 does the graph look something like this?